OpenStudy (anonymous):

the volume of a box is 192 cubic inches the height is 4 inches The width is 2 inches more than the length find the demintons desperate! i will give a medel

OpenStudy (anonymous):

volume = height*width*length = 192 height = 4 inches width = x length = x-2 (cuz width is 2 inches longer than length) volume = 192 = 4*x*(x-2) divide each sides by 4 48 = x(x-2) -> x(x-2)-48=0 -> x^2-2x-48=0 ->(x-8)(x+6)=0 x is either 8 or -6 but width cant be negative #, so the answer is height = 4 width = 8 length = 6